OpenStudy (megannicole51):

Find the sum 4+4(.5)+4(.5)^2+....+4(.5)^12 Sn=a1(1-r^n)/1-r n=13 a1=4 What's r?

OpenStudy (megannicole51):

@agent0smith

OpenStudy (agent0smith):

They make it easy by telling you r... which number is being multiplied each time?

OpenStudy (megannicole51):

.5?

OpenStudy (agent0smith):

Excellent :D now use the sum formula again

OpenStudy (megannicole51):

really thats r?!?! awesome!!!

OpenStudy (anonymous):

wow this one is way way (way) easier than the previous integrals, isn't it?

OpenStudy (agent0smith):

lol yes, maybe she's meant to be doing them via integration, but meh

OpenStudy (megannicole51):

nah this is easier! lol

OpenStudy (anonymous):

i hope not, but in any case this is a lot easier than finding the area in some sort of clover leaf using polar coordinates, or horrific partial fractions

OpenStudy (megannicole51):

so i got 7.999 and it says its wrong:(

OpenStudy (megannicole51):

(4(1-(.5)^13))/(1-(.5))

OpenStudy (anonymous):

i get the same thing

OpenStudy (megannicole51):

the question says whats the value of the sum

OpenStudy (megannicole51):

and yeah i double checked the question and its written out right

OpenStudy (megannicole51):

what am i doing wrong lol its probably something small!

OpenStudy (anonymous):

does it say "give whatevever decimal places " or "round to the nearest something"?

OpenStudy (agent0smith):

Hmmm :/

OpenStudy (megannicole51):

nope:/

OpenStudy (anonymous):

then stick with your answer, or write \(7.99902\)

OpenStudy (megannicole51):

its online homework:/

OpenStudy (megannicole51):

would i use a different equation and add 1 to the exponent? there was a formula my professor used and we did that but i have no idea what it was or why he was doing it lol....i really wish he spoke better english lol

OpenStudy (agent0smith):

I'm going to add them up one by one, on wolfram alpha.

OpenStudy (megannicole51):

do what u need to do:) lol

OpenStudy (anonymous):

i added, check link above

OpenStudy (agent0smith):

Yes but i meant type out each term :P

OpenStudy (anonymous):

you have more patience than i do

OpenStudy (anonymous):

oooh unless the question is really "what is r"? as written, in which case the answer is \(.5\) !!

OpenStudy (agent0smith):

copy and pasted +4(.5)^2 and edited the exponent a bunch of times, was quick

OpenStudy (anonymous):

is it one of those on line "step by step" questions, where they walk you through it? maybe the question only wants you to identify \(r\) after all, they have given you everything else, formula included

OpenStudy (megannicole51):

nope:/

OpenStudy (agent0smith):

@satellite73 I think Megan in all her genius is the one who gave everything after: "Find the sum 4+4(.5)+4(.5)^2+....+4(.5)^12"

OpenStudy (megannicole51):

its not .5 just btw i tried it:/

OpenStudy (anonymous):

ok then moveon.org you will not find another answer

OpenStudy (agent0smith):

Email your teacher @megannicole51 and tell her it's wrong (the online answer). If two guys with maths degrees (i'm assuming) and one university student can't figure it out, then it's obviously wrong ;)

OpenStudy (megannicole51):

he doesnt check his email until early mornings and this homework is due at 6am so its hopeless:/

OpenStudy (megannicole51):

hes a crazy old man bt hes so funny...i just wish he would check his email more

OpenStudy (agent0smith):

lol well yeah crazy old people don't be stayin on the innernet at night-time.

OpenStudy (megannicole51):

okay so the problem under it is.... 4(.5)^3+4(.5)^4+4(.5)^5+.....+4(.5)^11 so n=9 a1=4(.5)^3 so then what is r for this one? lets see if i can figure this one out and if we get this one wrong too then ill email my professor

OpenStudy (anonymous):

explain how \(n=9\)?

OpenStudy (agent0smith):

But this time it doesn't start at n=1. But, we can just subtract those terms from the sum...

OpenStudy (anonymous):

with props to @agent0smith

OpenStudy (megannicole51):

bc the square roots start at 3 and u just kinda count back...sorry thats a bad explanation lol

OpenStudy (megannicole51):

*square root lol that was super terrible grammar

OpenStudy (anonymous):

lol, they are exponents, not roots, square or otherwise

OpenStudy (megannicole51):

hahahahahah that just shows how tired i am...thank you

OpenStudy (agent0smith):

"super terrible grammar" is probably super terrible grammar ;)

OpenStudy (megannicole51):

thats why it was funny:) lol

OpenStudy (agent0smith):

:D

OpenStudy (megannicole51):

so any who.....how do we find r lol

OpenStudy (agent0smith):

4(.5)^3+4(.5)^4+4(.5)^5+.....+4(.5)^11 r looks the same as the last one

OpenStudy (ranga):

In a geometric series, 'r' is the common ratio and can be found by dividing any term by its previous term.

OpenStudy (megannicole51):

any 2 terms?

OpenStudy (ranga):

4(.5)^4 / 4(.5)^3 = 0.5 = r 4(.5)^5 / 4(.5)^4 = 0.5 = r

OpenStudy (ranga):

If you look at the series, the second term is the first term times .5 The third term is the second term times .5 So .5 is the common ratio.

OpenStudy (megannicole51):

awesome thank you:)

OpenStudy (ranga):

u r welcome!

OpenStudy (megannicole51):

okay so i got that second problem right! but i still cant figure out the first one!

OpenStudy (ranga):

Into Google I typed exactly "4(1 - .5^13)/(1 - .5) = " without the quotes and I get 7.9990234375.

OpenStudy (megannicole51):

@satellite73 hey so for the first problem i guess i just didnt put enough decimal points or something because i entered 7.9990234375 and it says it was correct! technology what are you guna do lol!

OpenStudy (megannicole51):

thank you i already got it:)

OpenStudy (agent0smith):

Haha awesome megan :)

OpenStudy (megannicole51):

thank you:D